OpenStudy (anonymous):

Differentiate: M(t)= [p/1-(1-p)e^t]^k

OpenStudy (turingtest):

\[M(t)=[\frac p{1-(1-p)e^t}]^k\]is the formula for M ?

OpenStudy (anonymous):

yep thats the formula i want to differentiate

OpenStudy (turingtest):

and p and k are constants, right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

Chain Rule\[(f\circ g)'(x)=f'(g(x))\cdot g'(x)\]In your case we can define the functions as so:\[f(x)=x^k\]\[g(x)=\frac p{1-(1-p)e^t}\]f can be done with the product rule, and g can be done with the quotient rule. Find each derivative and plug it into the product rule.

OpenStudy (anonymous):

ok thank you very much..so u positive this is the right method to go about doing this?

OpenStudy (turingtest):

Yes actually you can do the second part as\[g(x)=p[1-(1-p)e^t]^{-1}\]if you would rather use the chain rule again instead of the product rule, but you will get the same answer either way

OpenStudy (turingtest):

I meant instead of the quotient rule*

OpenStudy (anonymous):

ok il give it a try now

OpenStudy (turingtest):

Cool, let me know if you want me to post how I did it...

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

il be grateful

OpenStudy (turingtest):

\[M(t)=[\frac p{1-(1-p)e^t}]^k\]\[M'(t)=k[\frac p{1-(1-p)e^t}]^{k-1}\cdot(-p)[1-(1-p)e^t]^{-2}\cdot(p-1)e^t\]Now simplify...

OpenStudy (anonymous):

so u used the chain rule?

OpenStudy (turingtest):

yeah, twice

OpenStudy (turingtest):

\[f(t)=t^k\]\[g(t)=pt^{-1}\]\[h(t)=1-(1-p)e^t\]take each derivative and plug into the chain rule\[(f\circ g\circ h)'(t)=(f'\circ g\circ h)(t)\cdot(g'\circ h)(t)\cdot h'(t)\]sorry for changing t to x earlier, I hope that didn't confuse you

OpenStudy (anonymous):

so theres thhees functions

OpenStudy (anonymous):

three*

OpenStudy (turingtest):

there doesn't have to be, I just renamed the parts to make it easier you can also just say\[f(t)=t^k\]and\[g(t)=\frac p{1-(1-p)e^t}\]that way you only have two functions, but you will have to use the product or quotient rule on the second one

OpenStudy (anonymous):

Thank You very much..it seems that ur a pro at differentiation

OpenStudy (turingtest):

Differentiation is easy once you get used to it. Integration, on the other hand, remains tricky for much longer ;)

OpenStudy (anonymous):

sorry...how do u differentiate 1-(1-p)e^t

OpenStudy (turingtest):

the first one is a constant, so that becomes zero we can treat (1-p) as a constant coefficient, and d/dx(e^x)=e^x, so we get\[f(t)=1-(1-p)e^t\]\[f'(t)=-(1-p)e^t\]

OpenStudy (anonymous):

so the minus sign stays

OpenStudy (turingtest):

Yes, I think above I changed it like so:\[-(1-p)e^t=(p-1)e^t\]sorry if that threw you if you don't trust that analysis, you can distribute first\[1- (1-p)e^t=1-e^t+pe^t\]now take the derivative:\[-e^t+pe^t=(p-1)e^t\]so it works out the same either way, as it should

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)

OpenStudy (turingtest):

welcome :D

OpenStudy (anonymous):

after im done can u quickly check my answer for me please ?

OpenStudy (turingtest):

sure, but so can wolfram

OpenStudy (anonymous):

?

OpenStudy (turingtest):

There are a lot of different ways to simplify this answer, but mine came out looking like the first one under 'alternate forms' http://www.wolframalpha.com/input/?i=d%2Fdt%5Bp%2F%281-%281-p%29e%5Et%29%5D%5Ek

OpenStudy (turingtest):

If you are not familiar with wolfram, it is a good way to check answers. Don't take it as gospel though, I've caught making mistakes many times

OpenStudy (anonymous):

how would u simplify: [1-(1-p)e^t -p(p-1)e^t] / (1-(1-p)e^t)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

thats a mistake

OpenStudy (anonymous):

[1-(1-p)e^t -p(p-1)e^t] / (1-(1-p)e^t)^2

OpenStudy (anonymous):

how would u simplify the above

OpenStudy (turingtest):

I would change the numerator to\[1-(1-p)e^t +p(1-p)e^t\]and factor out (1-p)e^t:\[(1-1+p)(1-p)e^t=p(1-p)e^t\]and leave the denominator the same

OpenStudy (turingtest):

oh I think I messed up there XD on sec...

OpenStudy (anonymous):

cant u just get rid of: 1-(1-p)e^t from numerator and denominator?

OpenStudy (turingtest):

[1-(1-p)e^t -p(p-1)e^t]= [1-(1-p)e^t +p(1-p)e^t]=1-(1-p)(1-p)e^t but if that is your answer I think you are missing something... the way you have it written you could make it 1-[p(p-1)e^t] / (1-(1-p)e^t)^2 but it's hard to match up our answers because there are a lot of symbols I'll type my simplification...

OpenStudy (turingtest):

\[M'(t)=k[\frac p{1-(1-p)e^t}]^{k-1}\cdot(-p)[1-(1-p)e^t]^{-2}\cdot(p-1)e^t\]\[=\frac{pk(1-p)e^t[\frac p{1-(1-p)e^t}]^{k-1}}{[1-(1-p)e^t]^2}\]which is the same as wolfram's answer (the first one under 'Alternate forms')

OpenStudy (turingtest):

wolfram has a few +/- bits different than me, but you can reconcile that without too much difficulty through algebra they are the same

OpenStudy (turingtest):

...as far as how wolfram simplified to the answer at the top of the page, I don't feel like figuring that out. Seems excessive.

OpenStudy (anonymous):

no thats not my final answer

OpenStudy (anonymous):

i just need the simplification of what i gave above

OpenStudy (turingtest):

well without context it still looks wrong to me, but it is hard to tell I have to figure out what exactly you isolated in that part of the answer, and that's kinda hard to do from here...

OpenStudy (turingtest):

is that part you are asking about just what you got from the quotient rule?

OpenStudy (anonymous):

simplify: [1-(1-p)e^t + p(1-p)e^t]/ 1-(1-p)e^t)^2

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

yes from quotient rule

OpenStudy (turingtest):

I get, by factoring out (1-p)e^t\[\frac{1-(1-p)e^t + p(1-p)e^t}{ (1-(1-p)e^t)}^2=\frac{1+(1-p)(1-p)e^t}{1-(1-p)e^t}\]but I'm not so sure you did the quotient rule right, let me check

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

sorry the denomniator has a squared not the numerator

OpenStudy (anonymous):

so the denomaitor is: (1-(1-p)e^t)^2

OpenStudy (turingtest):

\[\frac p{1-(1-p)e^t}\]\[\frac{0-p[-(1-p)]e^t}{[1-(1-p)e^t]^2}=\frac{p(1-p)e^t}{[1-(1-p)e^t]^2}\]so I think you messed up on the derivative a bit

OpenStudy (turingtest):

...and that, times k[p/(1-(1-p)e^t)]^(k-1) gives my answer

OpenStudy (anonymous):

u=p, du/dp=1....v=1-(1-p)e^t, dv/dp=-(1-p)e^t..is that how i start?

OpenStudy (turingtest):

p is a constant though...

OpenStudy (turingtest):

dp/dt=0

OpenStudy (turingtest):

we are not taking the derivative w/respect to p

OpenStudy (turingtest):

u=p du/dt=0 v is correct though

OpenStudy (anonymous):

o silly me...sorry so it becomes zero lol?

OpenStudy (turingtest):

because p is constant I prefer to rewrite it as\[p[1-(1-p)e^t]^{-1}\]and use the chain rule instead. less messy in my opinion, but to each their own

OpenStudy (turingtest):

yes to your last question

OpenStudy (anonymous):

i dont know how to use chain rule with the above i get confused

OpenStudy (anonymous):

and v=-(1-p)e^t

OpenStudy (turingtest):

^that is dv/dt you mean here it is with the chain rule:\[p[1-(1-p)e^t]^{-1}=(-1)\cdot p[1-(1-p)e^t]^{-2}\cdot[-(1-p)e^t]\]\[=\frac{p(1-p)e^t}{[1-(1-p)e^t]^2}\]again, same answer

OpenStudy (anonymous):

thank you very much...sorry for my clumsiness

OpenStudy (turingtest):

Not at all, it's just a matter of practice ;) again, welcome!

OpenStudy (anonymous):

my final answer: k[p(1-p)e^t/[1-(1-p)e^t]^2] [p/1-(1-p)e^t]^k-1

OpenStudy (anonymous):

does this look promising?

OpenStudy (turingtest):

yes, I think that's right :D

OpenStudy (anonymous):

at last :D

OpenStudy (turingtest):

\[\frac{kp(1-p)e^t}{[1-(1-p)e^t]^2} \cdot[{1-(1-p)e^t}]^{k-1 }\]yeah that simplifies to my answer

OpenStudy (turingtest):

whoops dropped the p from the numerator, but you are right, don't worry

OpenStudy (anonymous):

LOL

OpenStudy (turingtest):

\[\frac{kp(1-p)e^t}{[1-(1-p)e^t]^2} \cdot[\frac p{1-(1-p)e^t}]^{k-1 }\]there we go, that's just a stone throw away from mine nicely done, congrads!

OpenStudy (anonymous):

Thank You...appreciate it

OpenStudy (anonymous):

sorry to have bothered u too much.

OpenStudy (turingtest):

Not at all, that's why I'm here

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

what methods would you use if we were to take the econd derivative of the answer we achieved above? so M''(t)?

OpenStudy (turingtest):

It's not going to be pretty... In that case it would be nice to get to wolframs factored form, then you could use the product rule. however with what we have, what I would probably do is write\[\frac{kp(1-p)e^t}{[1-(1-p)e^t]^2} \cdot[\frac p{1-(1-p)e^t}]^{k-1 }\]\[u=\frac{kp(1-p)e^t}{[1-(1-p)e^t]^2}\]\[v=[\frac p{1-(1-p)e^t}]^{k-1 }\]\[M''(t)=u'v+v'u\]no note that v' will be the same as the derivative we just found, but with k-1 instead of k. This way we can save a lot of work. That means you only need to find u', which can be done with the quotient rule. Note that you will need the chain rule on the denominator though, so yes, ugly answer expected.

OpenStudy (turingtest):

actually it's not as bad as I thought after simplification http://www.wolframalpha.com/input/?i=d%5E2%2Fdt%5E2%5Bp%2F%281-%281-p%29e%5Et%29%5D%5Ek I gotta eat, I'll check back later. Good luck

OpenStudy (turingtest):

also in finding u' you can take the kp(1-p) out (since it is constant) and just find the derivative of\[\frac{e^t}{[1-(1-p)e^t]^2}\]which should be a little less messy just don't forget to add back in kp(1-p) at the end